# Beyond Structure: Invariant Crystal Property Prediction with Pseudo-Particle Ray Diffraction

### v1

@@ 5 Experiments
We benchmark the models using three evaluation suites. The definitions and background of the regression and classification targets are introduced in the Appendix E.
  • Materials Project (MP) Database: We use stable structures retrieved from the Materials Project (Jain et al., 2020), comprising 122,959 entries with annotated formation energy, band gap, and metal/non-metal classification labels. Additionally, 9,473 of these entries include mechanical properties such as bulk modulus, shear modulus and Young’s modulus. The dataset is managed using the Atomic Simulation Environment (ASE), which provides a unified dataloader for streamlined access (Cao & Zhang, 2025b).
  • JARVIS-DFT (dft_3d): This dataset contains 75,993 entries, each annotated with formation energy, band gap, bulk modulus, and shear modulus (Choudhary et al., 2020).
@@ 5.1 Results
Method Form. Energy Bandgap Bulk Modulus Shear Modulus Young’s Modulus Metal/Non-metal
eV/atom eV log(GPa) log(GPa) log(GPa) Acc(%)
CGCNN (Xie & Grossman, 2018) 0.041 0.262 0.079 0.162 0.155 87.5
SchNet (Schütt et al., 2017) 0.038 0.344 0.092 0.129 0.143 82.8
MEGNET (Chen et al., 2019) 0.053 0.307 0.134 0.206 0.189 86.6
GATGNN (Louis et al., 2020) 0.083 0.343 0.051 0.111 0.109 86.3
Matformer (Yan et al., 2022) 0.035 0.193 0.089 0.306 0.291 92.7
Crystalformer (Taniai et al., ) 0.049 0.251 0.064 0.140 0.127 86.5
Crystalfarmer (Ito et al., 2025) 0.030 0.216 0.047 0.118 0.106 87.5
eComFormer (Yan et al., 2024) 0.033 0.153 0.088 0.269 0.311 82.5
PRDNet 0.028 0.151 0.035 0.108 0.104 93.3
Across the MP, JARVIS-DFT, and Matbench benchmarks, PRDNet consistently delivers superior or competitive performance compared with existing models. On the MP dataset (Table 1), PRDNet surpasses all baselines, achieving the lowest errors in formation energy (0.028 eV/atom), band gap (0.151 eV), bulk modulus (0.035 log(GPa)), shear modulus (0.108 log(GPa)), and Young’s modulus (0.104 log(GPa)), while also reaching the highest metal/non-metal classification accuracy (93.3%). On JARVIS-DFT (Table 2), PRDNet again outperforms all baselines, attaining the best results across all tasks, including formation energy (0.032 eV/atom), band gap (0.140 eV), bulk modulus (0.064 log(GPa)), and shear modulus (0.122 log(GPa)). On the Matbench benchmark (Table 2), PRDNet achieves the lowest errors in formation energy (0.019 eV/atom), shear modulus (0.058 log(GPa)) and refractive index (0.242 (unitless)), while maintaining competitive performance in exfoliation energy prediction. These results show that PRDNet generalizes effectively across diverse datasets and provides balanced improvements in both energy-related and mechanical property predictions, underscoring its robustness and versatility for crystal property modeling.
@@
Form. Energy Bandgap Bulk Modulus Shear Modulus jdft2d mp_e_form log_gvrh dielectric
eV/atom eV log(GPa) log(GPa) eV/atom eV/atom log(GPa) unitless
CGCNN 0.038 0.144 0.125 0.163 0.050 0.027 0.066 0.541
SchNet 0.074 0.278 0.143 0.225 0.056 0.042 0.085 0.341
MEGNET 0.074 0.221 0.124 0.214 0.052 0.036 0.098 0.459
GATGNN 0.069 0.235 0.078 0.160 0.046 0.032 0.087 0.408
Matformer 0.033 0.207 0.101 0.157 0.072 0.023 0.071 0.764
Crystalformer 0.052 0.248 0.075 0.146 0.071 0.020 0.080 0.244
Crystalframer 0.048 0.208 0.065 0.144 0.056 0.065 0.084 0.248
eComFormer 0.119 0.260 0.118 0.243 0.031 0.020 0.085 0.543
PRDNet 0.032 0.140 0.064 0.122 0.038 0.019 0.058 0.242
@@
Table 4: Pseudo-Particle v.s. X-ray Photon
Method MP JARVIS-DFT MB
Form. Energy (MAE eV/atom) 0.028 0.085 0.032 0.103 0.019 0.063
Bandgap (MAE eV) 0.151 0.271 0.140 0.331 - -
Bulk Modulus (MAE log(GPa)) 0.035 0.094 0.064 0.101 - -
Shear Modulus (MAE log(GPa)) 0.108 0.322 0.122 0.256 0.058 0.111
Young’s Modulus (MAE log(GPa)) 0.104 0.291 - - - -
Exfo. Energy (MAE eV/atom) - - - - 0.038 0.089
Refractive Index (MAE unitless) - - - - 0.242 0.476
Metal/Non-metal (Acc %) 93.3 78.9 - - - -
For evidencing the enriched representation of Pseudo-Particle on crystal sensitivity, we compare two PRDNet variants: (1) PRDNet-Learned: employs learned pseudo-particle form factors $f_{i}^{*}$ (Eq. 13); (2) PRDNet-Tabulated: substitutes $f_{i}^{*}$ with tabulated X-ray photon form factors. As summarized in Table 4, the pseudo-particle representation consistently outperforms the X-ray counterpart across all datasets and property prediction tasks. Specifically, it achieves lower MAEs for formation energy, bandgap, and mechanical moduli, while also delivering higher classification accuracy in metal/non-metal discrimination.
Number of Miller Indices
Each Miller index specifies a crystallographic direction. The choice of a sufficiently large set of Miller indices, $\mathcal{H}_{0}$, is crucial for adequately representing reciprocal space. We evaluate the performance of PRDNet using 100, 200, 300, 500, and 1000 Miller indices on the MP dataset for both formation energy and band gap prediction tasks. As shown in Figure 4, using 300 indices provides a good balance between computational cost and predictive accuracy.
@@ 6 Conclusion
In this work, we introduce a Pseudo-Particle Ray Diffraction Neural Network (PRDNet) for crystal structure property prediction. PRDNet outperforms previous approaches by capturing long-term atomic correlations through a physically invariant reciprocal-space representation embedded within the neural network. The learned pseudo-particle attributes act as a novel probe, exhibiting higher sensitivity than any known physical particle to atomic type, local chemical environment, and diffraction directions. Experiments demonstrate that PRDNet provides a principled framework for integrating crystallographic theory with deep learning, offering a novel conception of invariant crystal representation and achieving SOTA performance across multiple properties and databases.
Limitations
The benefits of PRDNet’s diffraction-based reciprocal-space representation do not directly transfer to non-periodic structures. Moreover, the current representation is difficult to invert for reconstructing structures from reciprocal-space sampling, limiting its applicability to direct crystal structure prediction tasks.
Future Work
Building on this work, future research may explore:
  • Extended Physics Integration: Incorporating additional physical phenomena, such as phonon interactions, electronic structures, and magnetic properties.
  • Multi-Scale Modeling: Expanding PRDNet to handle multiple length scales simultaneously, from atomic to mesoscopic structures.
  • Structure Identification: Integrating with experimental design frameworks for crystal structure identification (XQueryer, Cao et al. ), structure refinement (WPEM, Cao & Zhang (2024; 2025a)), and property inference (PRDNet) to accelerate materials discovery.
@@ Appendix A Additional Details
A.1 crystal invariance
@@
A.2 Physical Constraints of Diffraction
Recent work, such as ReGNet (Nie et al., 2025), incorporates reciprocal-space information to capture long-term atomic interactions. The model introduces a dedicated reciprocal block that approximates the structure factor, functioning in a manner similar to Eq. 3, and achieves improved performance. In ReGNet, reciprocal-space information is approximated through a simplified Fourier-like transformation. However, this approximation omits several critical components and consequently fails to preserve key physical properties inherent to reciprocal-space representations.
  • Dependence: The atomic form factor $f_{j}(\mathbf{Q})$, as defined in Eq. 3, is not a constant. It inherently depends on three fundamental factors: the atomic species, the local atomic environment, and the scattering vector $\mathbf{Q}$. Preserving these dependencies is crucial for maintaining physical fidelity in diffraction-based modeling.
  • Consistency: According to the Schrodinger equation(Berezin & Shubin, 2012), the charge density (or electron distribution) is uniquely determined by the atomic configuration. Therefore, the reciprocal-space representation must be consistent with this determinism. ReGNet introduces multiple reciprocal blocks that evolve independently, leading to inconsistencies and instability in the reciprocal embedding.
@@ A.2 Physical Constraints of Diffraction
where $h_{j}^{\ell}$ is a learnable embedding that evolves through network layers. This deviates from the physical structure factor $F(\mathbf{Q})$ by omitting key components:
  • Missing atomic form factor dependence: The diffraction factor $f_{j}(\mathbf{Q})$ should be expressed as:
    $f_{j}(\mathbf{Q})=f_{j}\big(|\mathbf{Q}|,G_{\theta}(\mathcal{G}),f_{j}^{\text{type}}\big)$ (26)
    where $|\mathbf{Q}|$ captures the dependence on the magnitude of the scattering vector, $G_{\theta}(\mathcal{G})$ reflects the local environment, and $f_{j}^{\text{type}}$ encodes atomic species. ReGNet’s embedding $h_{j}^{\ell}$ lacks this explicit structure.
  • Violation of reciprocal-space uniqueness: The Fourier transform $F(\mathbf{Q})$ of the charge density $\rho(\mathbf{r})$ is uniquely defined by a given atomic configuration. ReGNet’s use of multiple independently parameterized reciprocal blocks violates this uniqueness, potentially leading to non-physical or inconsistent representations.
A.3 Training Settings of PRDNET
A.3.1 Huber Loss for Outlier Robustness
@@
A.3.2 AdamW with Decoupled Weight Decay
@@
A.3.3 Cosine Annealing with Warm Restarts
@@
A.3.4 Layer Normalization with Learnable Parameters

### v2

@@ 5 Experiments
We benchmark the models on three standard evaluation suites, all independently retrieved from the official websites. The definitions and background of the regression and classification targets are introduced in the Appendix E.
  • Materials Project (MP) Database: We use stable structures retrieved from the Materials Project (Jain et al., 2020), comprising 122,959 entries with annotated formation energy, band gap, and metal/non-metal classification labels. Additionally, 9,473 of these entries include mechanical properties such as bulk modulus, shear modulus and Young’s modulus. The dataset is managed using the Atomic Simulation Environment (ASE), and is available at https://huggingface.co/datasets/caobin/CPPbenchmark.
  • JARVIS-DFT (dft_3d): This dataset contains 75,993 entries, each annotated with formation energy, band gap (calculated using either the OptB88vDW functional, denoted as OPT, or the TBMBJ functional, denoted as MBJ), bulk modulus, shear modulus, total energy (calculated using the OptB88vDW functional), and energy above the hull ($E_{\rm hull}$ ) (Choudhary et al., 2020). We also evaluate the baselines on the JARVIS-DFT-3D-2021 dataset (55,723 entries), which serves as an important supplementary benchmark.
@@ 5.1 Results
Method Param Num Form. Energy Bandgap Bulk Modulus Shear Modulus Young’s Modulus Metal/Non-metal
$M$ : million eV/atom eV log(GPa) log(GPa) log(GPa) Acc(%)
CGCNN (Xie and Grossman, 2018) 2.0MM 0.041 0.262 0.079 0.162 0.155 87.5
SchNet (Schütt et al., 2017) 0.3MM 0.038 0.344 0.092 0.129 0.143 82.8
MEGNET (Chen et al., 2019) 0.2MM 0.053 0.307 0.134 0.206 0.189 86.6
GATGNN (Louis et al., 2020) 0.6MM 0.083 0.343 0.051 0.111 0.109 86.3
Matformer (Yan et al., 2022) 15.4MM 0.035 0.193 0.089 0.306 0.291 92.7
Crystalformer (Taniai et al., ) 0.8MM 0.049 0.251 0.064 0.140 0.127 86.5
Crystalfarmer (Ito et al., 2025) 0.9MM 0.030 0.216 0.047 0.118 0.106 87.5
eComFormer (Yan et al., 2024) 5.6MM 0.033 0.153 0.088 0.269 0.311 82.5
EwaldMP (Kosmala et al., 2023) 12.2MM 0.052 0.332 0.098 0.175 0.186 84.2
PotNet (Lin et al., 2023) 1.8MM 0.035 0.251 0.104 0.175 0.151 88.9
ReGNet (Nie et al., 2025)  [^1] As of November 12, 2025, the ReGNet (ReciNet) code is unavailable; the results were obtained using our own implementation based on the paper’s settings. We use a hidden dimension of 304 instead of 256, as this brings the total parameter count close to the  3.3M reported in the paper; using 256 results in only 2.37M. 3.3MM 0.047 0.331 0.088 0.172 0.204 85.1
PRDNet 20.9MM 0.028 0.151 0.035 0.108 0.104 93.3
Across the MP, JARVIS-DFT, and Matbench benchmarks, PRDNet consistently delivers superior or competitive performance compared with existing models. On the MP dataset (Table 1), PRDNet surpasses all baselines, achieving the lowest errors in formation energy (0.028 eV/atom), band gap (0.151 eV), bulk modulus (0.035 log(GPa)), shear modulus (0.108 log(GPa)), and Young’s modulus (0.104 log(GPa)), while also reaching the highest metal/non-metal classification accuracy (93.3%). On JARVIS-DFT (Table 2), PRDNet again outperforms the baselines, achieving the best results on most tasks, including formation energy (0.032 eV/atom), band gap (OPT, 0.140 eV), bulk modulus (0.064 log(GPa)), shear modulus (0.122 log(GPa)), total energy (0.032 eV/atom) and band gap (MBJ, 0.267 eV) while also achieving competitive performance in $E_{\rm hull}$ (0.041 eV). On the Matbench benchmark (Table 2), PRDNet achieves the lowest errors in formation energy (0.019 eV/atom), shear modulus (0.058 log(GPa)) and refractive index (0.242 (unitless)), while maintaining competitive performance in exfoliation energy prediction. PRDNet also achieves SOTA performance on the JARVIS-DFT-3D-2021 dataset (except for the band gap (OPT) task) outperforming all baseline models (Table 8). These results show that PRDNet generalizes effectively across diverse datasets and provides balanced improvements in both energy-related and mechanical property predictions, underscoring its robustness and versatility for crystal property modeling.
@@
Form. Energy Bandgap Bulk Modulus Shear Modulus Total Energy Bandgap $E_{\rm hull}$ jdft2d mp_e_form log_gvrh dielectric
eV/atom (OPT) eV log(GPa) log(GPa) eV/atom (MBJ) eV eV eV/atom eV/atom log(GPa) unitless
CGCNN 0.038 0.144 0.125 0.163 0.041 0.402 0.049 0.050 0.027 0.066 0.541
SchNet 0.074 0.278 0.143 0.225 0.048 0.507 0.073 0.056 0.042 0.085 0.341
MEGNET 0.074 0.221 0.124 0.214 0.054 0.369 0.056 0.052 0.036 0.098 0.459
GATGNN 0.069 0.235 0.078 0.160 0.053 0.372 0.048 0.046 0.032 0.087 0.408
Matformer 0.033 0.207 0.101 0.157 0.041 0.341 0.034 0.072 0.023 0.071 0.764
Crystalformer 0.052 0.248 0.075 0.146 0.055 0.282 0.053 0.071 0.020 0.080 0.244
Crystalframer 0.048 0.208 0.065 0.144 0.047 0.270 0.050 0.056 0.065 0.084 0.248
eComFormer 0.119 0.260 0.118 0.243 0.035 0.335 0.034 0.031 0.020 0.085 0.543
EwaldMP 0.052 0.256 0.109 0.191 0.057 0.451 0.047 0.052 0.053 0.117 0.485
PotNet 0.034 0.144 0.112 0.156 0.039 0.339 0.079 0.033 0.029 0.084 0.397
ReGNet 1 0.054 0.201 0.092 0.178 0.041 0.355 0.051 0.045 0.033 0.079 0.512
PRDNet 0.032 0.140 0.064 0.122 0.032 0.267 0.041 0.038 0.019 0.058 0.242
@@
For evidencing the enriched representation of Pseudo-Particle on crystal sensitivity, we compare two PRDNet variants: (1) PRDNet-Learned: employs learned pseudo-particle form factors $f_{i}^{*}$ (Eq. 13); (2) PRDNet-Tabulated: substitutes $f_{i}^{*}$ with tabulated X-ray photon form factors. As summarized in Table 7 (Appendix B), the pseudo-particle representation consistently outperforms the X-ray counterpart across all datasets and property prediction tasks. Specifically, it achieves lower MAEs for formation energy, bandgap, and mechanical moduli, while also delivering higher classification accuracy in metal/non-metal discrimination.
Diffraction Module Assembling
Each real-space crystal representation baseline in Table 1 is integrated with our reciprocal-space diffraction module and the fusion mechanism defined in Eq.,7 and Eqs.,1417. This allows a systematic comparison of all baseline models after incorporating our diffraction module. Although several baselines are neither invariant nor equivariant, therefore cannot achieve the same level of physical consistency as PRDNet. We include them to provide a more comprehensive evaluation, as listed in Table 4. Overall, the results show that most models benefit from the unique reciprocal-space representation and the modality-level fusion mechanism. Exceptions include GATGNN and Crystalformer on shear modulus, Crystalformer on bulk modulus, and Matformer on metal/non-metal classification, which show decreased performance; all other models show improved or comparable results.
Table 4: Comparison of Mean Absolute Error and Accuracy on the MP dataset across baselines enhanced by diffraction modules. The models were trained on NVIDIA GeForce RTX 4090, GeForce RTX 4080, GeForce RTX 3090, NVIDIA A100 and A40 GPUs.
Method GPUs [^2] As these are supplemental experiments conducted within a limited time window, we were unable to train all models on the same GPUs. Therefore, we explicitly list the GPUs used for training each baseline for this table. Form. Energy Bandgap Bulk Modulus Shear Modulus Young’s Modulus Metal/Non-metal
eV/atom eV log(GPa) log(GPa) log(GPa) Acc(%)
[…]
MEGNET (Chen et al., 2019) $\times3$090 0.040 $\downarrow$ 0.291 $\downarrow$ 0.087 $\downarrow$ 0.202 $\downarrow$ 0.177 $\downarrow$ 87.1 $\uparrow$
[…]
@@ 6 Conclusion
In this work, we introduce a Pseudo-Particle Ray Diffraction Neural Network (PRDNet) for crystal structure property prediction. PRDNet outperforms previous approaches by capturing long-range atomic correlations through a physically invariant reciprocal-space representation embedded within the neural network. The learned pseudo-particle attributes act as a novel probe, exhibiting higher sensitivity than any known physical particle to atomic type, local chemical environment, and diffraction directions. Experiments demonstrate that PRDNet provides a principled framework for integrating crystallographic theory with deep learning, offering a novel conception of invariant crystal representation and achieving SOTA performance across multiple properties and databases. The limitation also arises from the fusion of complete reciprocal-space information, which implies that the inverse process for Crystal Structure Prediction (CSP), i.e., generating valid structures in real space by shaping distributions in diffraction space, can break invariance and thus cannot be directly applied to CSP. In future, we plan to: (1) Integrate PRDNet with the structure identification tool (XQueryer (Cao et al., 2025)) to enable a fully automated materials analysis workflow; (2) Extend PRDNet to simultaneously handle multiple length scales, from atomic to mesoscopic structures; (3) Extend relaxed invariance structures for CSP tasks.
Reproducibility statement
We are committed to ensuring the reproducibility of our research. To this end, we provide a publicly accessible GitHub repository accompanied by a comprehensive README file. The README provides detailed information on the PRDNet model, its hyperparameters, environment configuration, dependency installation, environment validation, datasets, training executor, and training procedures. Furthermore, Appendix D lists all benchmark model settings used for validation.
@@ Appendix A Additional Details
A.1 The Use of LLMs
In this work, we used LLMs as an auxiliary tool for polishing the writing and improving the clarity of the manuscript. Specifically, we use it to review the entire text for grammatical accuracy, improve sentence structure, and ensure consistent phrasing and tone throughout the paper.
A.2 crystal invariance
@@ A.1 The Use of LLMs
A.3 Physical Constraints of Diffraction
Recent works endeavor to introduce Fourier transforms to capture long-range interactions (Kosmala et al., 2023; Lin et al., 2023; Nie et al., 2025), incorporating reciprocal-space information to model long-range atomic correlations to some extent. However, a complete diffraction representation is uniquely invariant and physically determined; slight omissions or approximations may lead to incomplete information, often treated empirically as a form of supplemental atomic information fusion. In our work, we systematically address these dependencies and construct an framework that preserves invariance. Taking ReGNet (renamed ReciNet in the latest version) (Nie et al., 2025) as an example, the model introduces a dedicated reciprocal block that approximates the structure factor, functioning similarly to Eq. 3, and achieves improved performance. We further decompose their key Fourier-transform settings to discuss the underlying dependencies.
  • Dependence: The atomic form factor $f_{j}(\mathbf{Q})$, as defined in Eq. 3, is inherently depends on three fundamental factors: the atomic species, the local atomic environment, and the scattering vector $\mathbf{Q}$. Preserving these dependencies is crucial for maintaining physical fidelity in diffraction-based modeling.
  • Consistency: According to the Schrodinger equation(Berezin and Shubin, 2012), the charge density (or electron distribution) is uniquely determined by the atomic configuration. Therefore, the reciprocal-space representation must remain consistent with this determinism and cannot evolve independently.
@@ A.3 Physical Constraints of Diffraction
where $h_{j}^{\ell}$ is a learnable embedding that evolves through network layers. The following dependencies and invariances may be omitted in such formulations:
  • Form factor dependence: The complete diffraction factor $f_{j}(\mathbf{Q})$ should be expressed as
    $f_{j}(\mathbf{Q})=f_{j}\big(|\mathbf{Q}|,G_{\theta}(\mathcal{G}),f_{j}^{\text{type}}\big),$ (26)
    where $|\mathbf{Q}|$ captures the dependence on the magnitude of the scattering vector, $G_{\theta}(\mathcal{G})$ reflects the local atomic environment, and $f_{j}^{\text{type}}$ encodes the atomic species. The $h_{j}^{\ell}$ in Eq. 25 loses the dependence on $\mathbf{Q}$ and the local state $G_{\theta}(\mathcal{G})$, depending only on the atom and block orders.
  • Reciprocal-space uniqueness: The Fourier transform $F(\mathbf{Q})$ of the charge density $\rho(\mathbf{r})$ is uniquely defined by a given atomic configuration. Eq. 25 evolves the diffraction expression across blocks in a way that breaks this invariance, thereby losing the physical uniqueness.
  • Modal fusion: Diffraction captures long-range interactions by introducing complete reciprocal-space information, which reflects the periodic arrangement and physical interactions between the structure (i.e., electron attributes) and probing process. This reciprocal-space information is a holistic representation of the structure and thus should be fused at the modality level rather than at the atomic level.
A.4 Training Settings of PRDNET
A.4.1 Huber Loss for Outlier Robustness
@@
A.4.2 AdamW with Decoupled Weight Decay
@@
A.4.3 Cosine Annealing with Warm Restarts
@@
A.4.4 Layer Normalization with Learnable Parameters
@@ B.4 Edge Feature Ablation
Table 7: Pseudo-Particle v.s. X-ray Photon
Method MP JARVIS-DFT MB
Form. Energy (MAE eV/atom) 0.028 0.085 0.032 0.103 0.019 0.063
Bandgap (MAE eV) 0.151 0.271 0.140 0.331 - -
Bulk Modulus (MAE log(GPa)) 0.035 0.094 0.064 0.101 - -
Shear Modulus (MAE log(GPa)) 0.108 0.322 0.122 0.256 0.058 0.111
Young’s Modulus (MAE log(GPa)) 0.104 0.291 - - - -
Exfo. Energy (MAE eV/atom) - - - - 0.038 0.089
Refractive Index (MAE unitless) - - - - 0.242 0.476
Metal/Non-metal (Acc %) 93.3 78.9 - - - -
Number of Miller Indices : Each Miller index specifies a crystallographic direction. The choice of a sufficiently large set of Miller indices, $\mathcal{H}_{0}$, is crucial for adequately representing reciprocal space. We evaluate the performance of PRDNet using 100, 200, 300, 500, and 1000 Miller indices on the MP dataset for both formation energy and band gap prediction tasks. As shown in Figure 5, using 300 indices provides a good balance between computational cost and predictive accuracy.
@@ C.4 Step 4: Multi-Modal Fusion
C.5 Experiments on JARVIS-DFT-3D-2021
In Table 8, we present the performance of PRDNet and the baseline models on the JARVIS-DFT-3D-2021 dataset. All baseline models were independently retrained for a fair comparison. PRDNet consistently outperforms the baselines, achieving the lowest MAE in formation energy (0.025 eV/atom), bulk modulus (0.075 log(GPa)), shear modulus (0.132 log(GPa)), total energy (0.024 eV/atom), band gap (MBJ, 0.242 eV), and $E_{\rm hull}$ (0.040 eV). It also demonstrates competitive performance on the band gap (OPT) prediction task, with a MAE of 0.138 eV.
Furthermore, Table 9 summarizes the reused baseline MAE results reported in the respective original papers. PRDNet achieves the best performance on formation energy, total energy, and $E_{\rm hull}$, while ranking second on band gap (MBJ) and delivering competitive results on band gap (OPT).
Table 8: Mean Absolute Error comparison on the JARVIS-DFT-3D-2021 dataset. We adopt the same data-splitting strategy as CrystalFormer (Taniai et al., ), applying a random seed of 123 and dividing the dataset into training, validation, and test sets with an 8:1:1 ratio.
Method Form. Energy Bandgap Bulk Modulus Shear Modulus Total Energy Bandgap $E_{\rm hull}$
eV/atom (OPT) eV log(GPa) log(GPa) eV/atom (MBJ) eV eV
CGCNN 0.039 0.164 0.107 0.153 0.044 0.374 0.099
SchNet 0.048 0.219 0.109 0.162 0.067 0.459 0.198
MEGNET 0.045 0.175 0.120 0.201 0.048 0.409 0.089
GATGNN 0.042 0.201 0.107 0.155 0.051 0.450 0.079
Matformer 0.041 0.166 0.334 0.386 0.041 0.311 0.057
Crystalformer 0.032 0.125 0.077 0.141 0.034 0.305 0.060
Crystalframer 0.038 0.132 0.081 0.134 0.042 0.330 0.070
eComFormer 0.037 0.165 0.200 0.275 0.035 0.301 0.055
EwaldMP 0.059 0.209 0.120 0.169 0.084 0.451 0.120
PotNet 0.034 0.141 0.106 0.155 0.062 0.301 0.079
ReGNet 0.038 0.195 0.211 0.162 0.047 0.336 0.071
PRDNet 0.025 0.138 0.075 0.132 0.024 0.242 0.040
Table 9: Mean Absolute Error comparison on the JARVIS-DFT-3D-2021 dataset. We adopt the baseline results reported by ReGNet (Nie et al., 2025) and Crystalframer (Ito et al., 2025).
Method Form. Energy Bandgap Total Energy Bandgap $E_{\rm hull}$
eV/atom (OPT) eV eV/atom (MBJ) eV eV
CGCNN 0.063 0.200 0.078 0.410 0.170
SchNet 0.045 0.190 0.047 0.430 0.140
MEGNET 0.047 0.145 0.058 0.340 0.084
GATGNN 0.047 0.170 0.056 0.510 0.120
Matformer 0.033 0.137 0.035 0.300 0.064
Crystalformer 0.031 0.128 0.032 0.274 0.046
Crystalframer 0.026 0.117 0.028 0.242 0.047
eComFormer 0.028 0.124 0.032 0.280 0.044
PotNet 0.029 0.127 0.032 0.270 0.055
ReGNet 0.027 0.126 0.027 0.240 0.043
PRDNet 0.025 0.138 0.024 0.242 0.040
C.6 Experiments on Materials Project (MEGNet) dataset
In Table 10, we summarize the performance of PRDNet and the baseline models on the Materials Project (MEGNet) dataset. PRDNet achieves the lowest MAE for bulk modulus (0.031 log(GPa)) and shear modulus (0.062 log(GPa)), and a competitive result for band gap (0.187 eV).
Table 10: Mean Absolute Error comparison on the Materials Project (MEGNet) dataset. We adopt the baseline results reported by ReGNet (Nie et al., 2025).
Method Form. Energy Bandgap Bulk Modulus Shear Modulus
eV/atom eV log(GPa) log(GPa)
CGCNN 0.031 0.292 0.047 0.077
SchNet 0.033 0.345 0.066 0.099
MEGNET 0.030 0.307 0.060 0.099
GATGNN 0.033 0.280 0.045 0.075
Matformer 0.021 0.211 0.043 0.073
Crystalformer 0.019 0.198 0.038 0.069
Crystalframer 0.017 0.185 0.034 0.068
eComFormer 0.018 0.202 0.042 0.073
PotNet 0.019 0.204 0.040 0.065
ReGNet 0.017 0.189 0.033 0.063
PRDNet 0.020 0.187 0.031 0.062